# Valid F-screening in linear regression

### v1

@@ 6.2 Retrospective inference
Keil et al. (2023) explores the relationship between sleep and cognitive impairment in older adults. The authors considered four response variables: (i) Center for Epidemiologic Studies-Depression Scale (CES-D), (ii) sleep duration, (iii) Mini-Mental State Examination (MMSE), and (iv) Mattis Dementia Rating Scale (MDRS) (see references in Keil et al. (2023) for details on these measures). For each response variable ($Y$) corresponding to the rows in Fig. 1, Keil et al. (2023) fit a model of the form
$Y=\beta_{1}\cdot\bm{I}_{<65}+\beta_{2}\cdot\bm{I}_{\text{65--85}}+\beta_{3}\cdot\bm{I}_{>85}+\epsilon,$
where $\bm{I}_{<65}$, $\bm{I}_{\text{65--85}}$, and $\bm{I}_{>85}$ denote indicator variables for the age group (in years) of $n=826$ study participants. The second-to-last column of Fig. 1 displays the p-values for a test of $H_{0}^{1:p}:\beta_{1}=\beta_{2}=\beta_{3}=0$ using ${F_{H_{0}^{1:p}}}$ from (3). The last column of Fig. 1 reports p-values for standard tests of $H_{0}^{j}:\beta_{j}=0$ for $j=1,2,3$.
@@
Keil et al. (2023) are not explicit about whether they would have tested $H_{0}^{j}$, $j=1,2,3$ had they not rejected $H_{0}^{1:p}$, and it is not possible to ascertain their intentions from their results, since in their Table 1 (a subset of which is displayed in our Fig. 1), all hypotheses of the form $H_{0}^{1:p}$ tested with ANOVA are rejected. Notably, they refer to tests of $H_{0}^{j}$ as “post hoc tests”; this phrasing is typically reserved in the applied literature to refer to tests that are conducted after rejecting $H_{0}^{1:p}$ (Everitt & Skrondal, 2010). Thus, for the sake of demonstration we shall make the not unreasonable assumption that they would not have tested $H_{0}^{j}$ for $j=1,2,3$ in their Table 1 if $H_{0}^{1:p}$ had not been rejected.
@@
Thus, we apply the methods in Section 3 and 4 to conduct a retrospective analysis, using only the summary statistics reported in Fig. 1, i.e. without access to the raw data. For each of the four outcomes in Fig. 1, we compute ${p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}}$ in (17) for $H_{0}^{j}:\beta_{j}=0$, for $j=1,2,3$, alongside the standard p-value in (5) which does not account for the fact that $H_{0}^{1:p}$ was rejected. The standard p-values that we compute match those in Fig. 1 if the Šidák correction is applied (Šidák, 1967).
@@
The results are displayed in Table 2. Generally, the values of ${p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}}$, as defined in (17), are equal to or larger than the corresponding values of $p_{H_{0}^{M}}$ as given by (5). For example, the p-value for the mean difference in CES-D between the under-65 age group and the 65-85 age group is 0.003 using $p_{H_{0}^{M}}$ versus 0.117 using ${p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}}$, and the p-value for the mean difference in CES-D between the 65-85 age group and over-85 age group is 0.016 versus 0.099. At the significance level of 0.05 used by Keil et al. (2023), the selective p-value ${p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}}$ does not lead to rejection for those two post hoc tests, whereas the standard pvalue $p_{H_{0}^{M}}$ does.
@@ 7 Discussion
We have shown that a common practice in applied statistics, of reporting the outputs of a linear regression model if and only if an overall $F$-test is rejected, leads to invalid inference. We have proposed an easily-implemented approach, based on conditional selective inference, that remedies the problem. We present a toolbox for valid inference in this setting that contains the most critical linear model outputs: p-values, confidence intervals, and point estimates. Furthermore, our selective p-values can be computed even without access to the raw data, using only the standard outputs of a multiple linear regression.
@@
  • Our results relied on an assumption that the errors are Gaussian. Can this be relaxed beyond Gaussianity, using ideas from e.g. Tian & Taylor (2017); Panigrahi (2023)? (We note, of course, that standard inference for a linear model also assumes Gaussianity in finite samples—thus, reliance on Gaussianity is only a very minor limitation.)
@@
  • How should our selective p-values be adjusted to account for multiple comparisons? For instance, how can ${p_{H_{0}^{M}\mid E}}$ for $M\in\{M_{1},\ldots,M_{L}\}$ be corrected in order to control family-wise error rate? An extension of Scheffé’s correction (Scheffé, 1953) to the selective setting may be possible.
@@ Acknowledgements
The authors gratefully acknowledge funding from NSF DMS 2322920, NIH 5P30DA048736, and ONR N00014-23-1-2589 to DW; NSF DMS-2303371, the Pacific Institutefor the Mathematical Sciences, the Simons Foundation, and the eScience Institute at the University of Washington to DK.
Appendix A Multiple testing correction is not sufficient for valid F-Screening
It is natural to wonder whether the problem addressed in this paper could be more simply tackled using a multiplicity correction for the number of post hoc tests performed in Step 2 of Box 1. In fact, as discussed in Section 1, applying a multiplicity correction to $p$ hypothesis tests in Step 2 of Box 1 will not, in general, yield Type 1 error control of any individual test, conditional on the fact that $H_{0}^{1:p}$ was rejected. We see this in the right-hand panel of Fig. A.1.
@@ Appendix A Multiple testing correction is not sufficient for valid F-Screening
Figure A.1: We simulate data according to (1) with $n=100$, $p=3$, $\sigma^{2}=1$, and $\beta_{1}=\beta_{2}=\beta_{3}=0$, so that $H_{0}^{1:p}$ (2) holds. For the test of $H_{0}^{1}:\beta_{1}=0$, we report (i) the uncorrected p-value that does not account for F-screening, i.e. $p_{H_{0}^{M}}$ in (5); (ii) $2p_{H_{0}^{M}}$, resulting from a Bonferroni correction to $p_{H_{0}^{M}}$ that accounts for the fact that two tests ($H_{0}^{1:p}$ and $H_{0}^{1}$) were conducted; (iii) $3p_{H_{0}^{M}}$, a Bonferroni correction that accounts for the fact that three post hoc tests ($H_{0}^{j}:\beta_{j}=0$ for $j\in\left\{1,2,3\right\}$) could have been conducted; (iv) the p-value that results from using Scheffé’s method; and (v) the selective p-value (12) that accounts for F-screening and uses the true value of error variance. Left: For each of 100,000 simulated datasets, the QQ-plot displays the p-values for $H_{0}^{1}$. The uncorrected p-value $p_{H_{0}^{M}}$ follows a $\operatorname{Unif}(0,1)$ distribution, while the selective p-values are extremely conservative because all results for all simulated datasets are displayed, regardless of rejection of $H_{0}^{1:p}$. Since only a single test has been conducted, the Bonferroni-corrected and Scheffé-corrected p-values are also conservative. Right: For the 5% of the 100,000 simulated datasets for which $H_{0}^{1:p}$ is rejected at level $\alpha_{0}=0.05$, we display the p-values for $H_{0}^{1}$. The uncorrected p-value $p_{H_{0}^{M}}$ is anti-conservative, as are the Bonferroni-corrected and Scheffé-corrected p-values when small quantiles are considered. However, the selective p-value follows a $\operatorname{Unif}(0,1)$ distribution.
@@ B.1 Point estimates
Then, the conditional density of $Y$ given $Y\in E_{1}$ and $Y\in\tilde{E}(\tilde{y})$ is
$f_{Y\mid Y\in E_{1}\cap\tilde{E}(\tilde{y})}(y)=\frac{f_{\tilde{Z}}\left(D_{1}y\right)f_{\tilde{W}}\left(D_{2}y\right)\bm{I}\left\{y^{\top}D_{2}y-y^{\top}D_{1}y\geq-\tilde{d}^{\top}\tilde{d}\right\}\bm{I}\left\{D_{3}y=\tilde{d}\right\}}{\Pr\left(\tilde{W}^{\top}\tilde{W}-c\tilde{Z}^{\top}\tilde{Z}\geq\tilde{d}^{\top}\tilde{d}\right).},$ (B.1)
where the marginal densities of $\tilde{Z}$ and $\tilde{W}$ are given in the proof of Lemma E.1.
@@
First, we can characterize $E_{1}$ in terms of $\left(\tilde{Z},\tilde{W},\tilde{V}\right)$ by observing that (almost surely)
@@
where the third implication exploits the idempotence and symmetry of projection matrices. We can similarly characterize $\tilde{E}(\tilde{y})$ in terms of $\tilde{V}$ by noting that
$Y\in\tilde{E}(\tilde{y})$ $\iff P_{X_{-j}}Y=P_{X_{-j}}\tilde{y}$ (B.3)
@@
Taken together, (B.2) and (B.3) enable us to characterize $E_{1}\cap\tilde{E}(\tilde{y})$ in terms of $\left(\tilde{Z},\tilde{W},\tilde{V}\right)$ by observing that (almost surely)
@@
Next, note that because the transformation from $Y$ to $\left(\tilde{Z},\tilde{W},\tilde{V}\right)$ is given by $D$, and $D^{\top}D=I_{n-1}$, a change of variables is straightforward:
@@
Then, we obtain the conditional density of $Y$ given $Y\in\tilde{E}(\tilde{y})$ and find
@@
$=f_{\tilde{Z}}\left(D_{1}y\right)f_{\tilde{W}}\left(D_{2}y\right)\bm{I}\left\{D_{3}y=\tilde{d}\right\},$
@@
Next, we (additionally) condition on $Y\in E_{1}$ (since we had already conditioned on $Y\in\tilde{E}(\tilde{y})$ in (B.6), we are now effectively conditioning on $Y\in E_{1}\cap\tilde{E}(\tilde{y})$). Standard application of the definition of conditional density yields
@@
$=\frac{f_{Y\mid\tilde{V}=\tilde{d}}(y)\bm{I}\left\{y^{\top}D_{2}y-cy^{\top}D_{1}y\geq-\tilde{d}^{\top}\tilde{d}\right\}}{\Pr\left(\tilde{W}^{\top}\tilde{W}-c\tilde{Z}^{\top}\tilde{Z}\geq-\tilde{d}^{\top}\tilde{d}\right)}$
$=\frac{f_{\tilde{Z}}\left(D_{1}y\right)f_{\tilde{W}}\left(D_{2}y\right)\bm{I}\left\{D_{3}y=\tilde{d}\right\}\bm{I}\left\{y^{\top}D_{2}y-cy^{\top}D_{1}y\geq-\tilde{d}^{\top}\tilde{d}\right\}}{\Pr\left(\tilde{W}^{\top}\tilde{W}-c\tilde{Z}^{\top}\tilde{Z}\geq-\tilde{d}^{\top}\tilde{d}\right)},$
@@
We now comment on Fig. B.1. When the overall null $H_{0}^{1:p}$ (2) holds, the standard MLE has a much larger spread than the conditional MLE derived in this section, and indeed the mode of the former is not even at 00. This is because the overall null is rejected when the F-statistic is sufficiently large, which tends to occur when $\hat{\beta}_{1}$ and/or $\hat{\beta}_{2}$ is large. Thus, the standard MLE for $\beta_{1}$ will tend to be far from zero conditional on the rejection of the overall null. The selective estimate, on the other hand, adjusts for the selection event, and therefore has its mode at the true value 00 and is more tightly concentrated around it.
[…]

### v2

@@ 6.2 Retrospective inference
Keil et al. (2023) explores the relationship between sleep and cognitive impairment in older adults. The authors considered four response variables: (i) Center for Epidemiologic Studies-Depression Scale (CES-D), (ii) sleep duration, (iii) Mini-Mental State Examination (MMSE), and (iv) Mattis Dementia Rating Scale (MDRS) (see references in Keil et al. (2023) for details). For each response variable corresponding to the rows in Fig. 1, Keil et al. (2023) fit a model of the form $Y=\beta_{1}\cdot\bm{I}_{<65}+\beta_{2}\cdot\bm{I}_{\text{65--85}}+\beta_{3}\cdot\bm{I}_{>85}+\epsilon,$ where $\bm{I}_{<65}$, $\bm{I}_{\text{65--85}}$, and $\bm{I}_{>85}$ denote indicator variables for the age group (in years) of $n=826$ study participants. The second-to-last column of Fig. 1 displays the p-values for a test of $H_{0}^{1:p}:\beta_{1}=\beta_{2}=\beta_{3}=0$, and the last column reports p-values for standard tests of $H_{0}^{j}:\beta_{j}=0$ for $j=1,2,3$.
@@
Keil et al. (2023) are not explicit about whether they would have tested $H_{0}^{j}$, $j=1,2,3$ had they not rejected $H_{0}^{1:p}$, and it is not possible to ascertain their intentions from their results, since in their Table 1 (a subset of which is displayed in our Fig. 1), all hypotheses of the form $H_{0}^{1:p}$ tested with ANOVA are rejected. Notably, they refer to tests of $H_{0}^{j}$ as “post hoc tests”; this phrasing is typically reserved in the applied literature to refer to tests that are conducted after rejecting $H_{0}^{1:p}$ (Everitt & Skrondal, 2010). Thus, for the sake of demonstration we shall assume that they would not have tested $H_{0}^{j}$ for $j=1,2,3$ in their Table 1 if $H_{0}^{1:p}$ had not been rejected.
@@
Accordingly, we apply the methods in Section 3 and 4 to conduct a retrospective analysis, using only the summary statistics reported in Fig. 1. For each of the four outcomes in Fig. 1, we compute $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$} in (21) for $H_{0}^{j}:\beta_{j}=0$, for $j=1,2,3$, alongside the standard p-value in (5) which does not account for the fact that $H_{0}^{1:p}$ was rejected. The standard p-values that we compute match those in Fig. 1 if the Šidák correction is applied (Šidák, 1967).
@@
The results are displayed in Table 2. Generally, the values of $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$} (21) are equal to or larger than the corresponding values of $p_{H_{0}^{M}}$ (5). For example, the p-value for the mean difference in CES-D between the under-65 age group and the 65-85 age group is 0.003 using $p_{H_{0}^{M}}$ versus 0.117 using $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$}, and the p-value for the mean difference in CES-D between the 65-85 age group and over-85 age group is 0.016 versus 0.099. At the significance level of 0.05 used by Keil et al. (2023), the selective p-value $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$} does not lead to rejection for those two post hoc tests, whereas the standard p-value $p_{H_{0}^{M}}$ does.
@@ 7 Discussion
We have shown that a common practice in applied statistics of reporting the outputs of a linear regression model if and only if an overall $F$-test is rejected leads to invalid inference. We have proposed an easily-implemented approach, based on conditional selective inference, that remedies the problem. We present a toolbox for valid inference in this setting that contains the most critical linear model outputs: p-values, confidence intervals, and point estimates. Furthermore, our selective p-values can be computed even without access to the raw data, using only the standard outputs of a multiple linear regression.
@@
  • Our results relied on an assumption that the errors are Gaussian. Can this be relaxed using ideas from e.g. Tian & Taylor (2017); Panigrahi (2023)? (We note, of course, that standard inference for a linear model also assumes Gaussianity in finite samples—thus, reliance on Gaussianity is only a very minor limitation of our method.)
@@
  • How should our selective p-values be adjusted to account for multiple comparisons? For instance, how can $p_{H_{0}^{M}\mid E}$} for $M\in\{M_{1},\ldots,M_{L}\}$ be corrected in order to control family-wise selective Type 1 error rate? An extension of Scheffé’s correction (Scheffé, 1953) to the selective setting may be possible.
@@ Acknowledgements
The authors gratefully acknowledge funding from NSF DMS 2322920, NSF DMS 2514344, NIH 5P30DA048736, and ONR N00014-23-1-2589 to DW; NSF DMS-2303371, the Pacific Institute for the Mathematical Sciences, the Simons Foundation, and the eScience Institute at the University of Washington to DK; and NSF Graduate Research Fellowship Grant No. DGE-2140004 to OM.
Appendix A Connections to multiple testing
A.1 Family-wise error rate control does not imply selective Type 1 error rate control
One might wonder whether the problem addressed in this paper could be more simply tackled by controlling the family-wise error rate (FWER). FWER control has been studied for almost a century (Tukey, 1949; Scheffé, 1953; Dunn, 1961; Holm, 1979; Marcus et al., 1976; Hochberg & Tamhane, 1987), and continues to be an area of active research: for instance, the “PoSI” proposal of Berk et al. (2013)—foundational to the field of selective inference—targets the FWER.
However, we see in the right-hand panel of Fig. A.1 that, in the context of F-screening in a setting with $p=3$, control of the FWER does not imply control of the selective Type 1 error rate. In particular, though Bonferroni, Scheffé, and closed testing corrections are guaranteed to control FWER, they do not control the Type 1 error rate conditional on rejection of $H_{0}^{1:3}$.
@@ A.1 Family-wise error rate control does not imply selective Type 1 error rate control
Figure A.1: We simulate data according to (1) with $n=100$, $p=3$, $\sigma^{2}=1$, and $\beta_{1}=\beta_{2}=\beta_{3}=0$, so that $H_{0}^{1:p}$ (2) holds. For the test of $H_{0}^{1}:\beta_{1}=0$, we report (i) the uncorrected p-value that does not account for F-screening, i.e. $p_{H_{0}^{M}}$ in (5); (ii) $2p_{H_{0}^{M}}$, resulting from a Bonferroni correction to $p_{H_{0}^{M}}$ that accounts for the fact that two tests ($H_{0}^{1:p}$ and $H_{0}^{1}$) were conducted; (iii) $3p_{H_{0}^{M}}$, a Bonferroni correction that accounts for the fact that three post hoc tests ($H_{0}^{j}:\beta_{j}=0$ for $j\in\left\{1,2,3\right\}$) could have been conducted; (iv) the p-value that results from using Scheffé’s method; (v) the p-value associated with the closed testing procedure as described in Section A.2; and (vi) the selective p-value (10) that accounts for F-screening and uses the true value of error variance. Left: For each of 100,000 simulated datasets, the $Q$-plot displays the p-values for $H_{0}^{1}$. The uncorrected p-value $p_{H_{0}^{M}}$ follows a $\operatorname{Unif}(0,1)$ distribution, while the selective p-values are extremely conservative because all results for all simulated datasets are displayed, regardless of rejection of $H_{0}^{1:p}$. Since we are only considering a single test, the Bonferroni-corrected, Scheffé-corrected, and closed-testing p-values are also conservative. Right: For the 5% of the 100,000 simulated datasets for which $H_{0}^{1:p}$ is rejected at level $\alpha_{0}=0.05$, we display the p-values for $H_{0}^{1}$. The uncorrected p-value $p_{H_{0}^{M}}$ and the closed testing p-value are anti-conservative, as are the Bonferroni-corrected and Scheffé-corrected p-values when small quantiles are considered. However, the selective p-value follows a $\operatorname{Unif}(0,1)$ distribution.
Why is this the case? Consider the setting of Fig. A.1. There are three possibilities associated with the null hypotheses $H_{0}^{1:3}:\beta_{1:3}=0$ and $H_{0}^{1}:\beta_{1}=0$: (i) both are true; (ii) $H_{0}^{1:3}$ is false, but $H_{0}^{1}$ is true; or (iii) both are false. In (iii), a Type 1 error is impossible, so we consider only (i) and (ii). Let $\delta_{1:3}$ be the decision function associated with the test of $H_{0}^{1:3}$: it is a function of the (random) data and takes value $1$ when we reject $H_{0}^{1:3}$ and value 0 otherwise; let $\delta_{1}$ similarly denote the decision function associated with testing $H_{0}^{1}$. In scenario (i), the FWER is ${\mathbb{E}}_{H_{0}^{1:3}}\left[\max\{\delta_{1:3},\delta_{1}\}\right]$. In scenario (ii), the FWER is ${\mathbb{E}}_{H_{0}^{1}}\left[\delta_{1}\right]$. By contrast, in scenario (i) the selective Type 1 error — which is controlled by our procedure — is ${\mathbb{E}}_{H_{0}^{1:3}}\left[\delta_{1}\mid\delta_{1:3}=1\right]$, while in scenario (ii) it is ${\mathbb{E}}_{H_{0}^{1}}\left[\delta_{1}\mid\delta_{1:3}=1\right]$. Clearly, the former targets are not the same as the latter targets, and control of one does not imply control of the other.
A.2 Connection to closed testing
At first glance, there appears to be a connection between the F-screening procedure described in Box 1, and so-called “closed testing procedures” (Marcus et al., 1976).
We begin by describing the closed testing procedure of Marcus et al. (1976), which controls the FWER. Let $H_{1},H_{2},\dots,H_{q}$ denote the set of elementary hypotheses to be tested. In addition to the elementary hypotheses, we also consider all possible intersections $H_{I}=\cap_{i\in I}H_{i}$ for all nonempty $I\subseteq\left\{1,2,\dots,q\right\}$. Next, we record the results of a (preliminary) test at level $\alpha$ for each of the hypotheses (including the elementary and intersection hypotheses). We then consider an elementary hypothesis $H_{i}$ to be rejected if and only if the preliminary test $H_{I}$ is rejected for all $H_{I}$ where $i\in I$.
How can we apply closed testing in our setting? First, consider elementary hypotheses $H_{0}^{M}:\beta_{M}=0$ and $H_{0}^{-M}:\beta_{-M}=0$. In order to conduct a closed testing procedure, the only additional hypothesis we need to consider is $H_{0}^{1:p}=H_{0}^{M}\cap H_{0}^{-M}$. In order to reject $H_{0}^{M}$, closed testing dictates that our preliminary tests must have rejected $H_{0}^{1:p}$ and $H_{0}^{M}$; of course, we can conduct the preliminary test of the former before we conduct the preliminary test of the latter. Thus, in this setting, closed testing bears a resemblance to the procedure in Box 1. The p-value associated with the test of $H_{0}^{M}$ in this particular instance of closed-testing is the maximum of the two p-values associated with the preliminary tests of $H_{0}^{1:p}$ and $H_{0}^{M}$. In Fig. A.1, we plot this p-value in the case that $p=3$ and the test of interest in Step 2 of Box 1 is $H_{0}^{1}:\beta_{1}=0$. We see that the closed testing p-value does not control the selective Type 1 error.
How can we reconcile the FWER guarantees of closed testing with the empirical results depicted in the right-hand panel of Figure A.1? The issue hinges on how we “count” the decision regarding $H_{0}^{M}$ in scenarios wherein we fail to reject $H_{0}^{1:p}$. We illustrate this point with a simple example, in which $H_{0}^{1:p}$ holds. Suppose that one hundred researchers independently collect data and then perform closed testing at level $\alpha$. This entails conducting preliminary tests of $H_{0}^{1:p}$, $H_{0}^{M}$, and $H_{0}^{-M}$. Suppose that five of the researchers’ preliminary tests reject $H_{0}^{1:p}$; of these five, four of their preliminary tests also reject $H_{0}^{M}$. Then, the closed testing procedure dictates that these four researchers should reject $H_{0}^{M}$. In this setting, the family-wise error proportion is $5/100=5\%$, because five of the 100 researchers falsely rejected at least one null hypothesis: this is consistent with the fact that closed testing guarantees FWER control. In contrast, the resulting selective Type 1 error proportion — defined as the proportion of researchers who rejected $H_{0}^{M}$ among those who rejected $H_{0}^{1:p}$ — is $4/5=80\%$; this is because closed testing does not control the selective Type 1 error rate.
While other work on closed testing has built upon the proposal of Marcus et al. (1976) and has considered other targets besides FWER (Goeman & Solari, 2011), this literature does not target the selective Type 1 error rate that is the focus of our paper.
@@ B.1 Point estimates
Define new random variables $\tilde{Z}=D_{1}Y$, $\tilde{W}=D_{2}Y$, and $\tilde{V}=D_{3}Y$ (note that $\tilde{Z}^{\top}\tilde{Z}/\sigma^{2}\overset{d}{=}Z$ and $\tilde{W}^{\top}\tilde{W}/\sigma^{2}\overset{d}{=}W$, for $Z$ and $W$ that appear in Proposition 2.1). Let $c$ be as defined in (11) and $E_{1}$ as defined in (6), let $\tilde{y}$ denote a realization of $Y$, and define $\tilde{d}=P_{X_{-j}}\tilde{y}$. Define $\tilde{E}(\tilde{y})=\left\{Y:P_{X_{-j}}Y=P_{X_{-j}}\tilde{y}\right\}$.
@@
$f_{Y\mid Y\in E_{1}\cap\tilde{E}(\tilde{y})}(y)=\frac{f_{\tilde{Z}}\left(D_{1}y\right)f_{\tilde{W}}\left(D_{2}y\right)\bm{I}_{\left\{y^{\top}D_{2}y-y^{\top}D_{1}y\geq-\tilde{d}^{\top}\tilde{d}\right\}}(y)\bm{I}_{\left\{D_{3}y=\tilde{d}\right\}}(y)}{\Pr\left(\tilde{W}^{\top}\tilde{W}-c\tilde{Z}^{\top}\tilde{Z}\geq\tilde{d}^{\top}\tilde{d}\right).}$, (B.1)
where the marginal densities of $\tilde{Z}$ and $\tilde{W}$ are given in (B.8) and $\bm{I}_{y}\{\cdot\}$ denotes the indicator that $y$ satifies the condition inside the brackets.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
$=f_{\tilde{Z}}\left(D_{1}y\right)f_{\tilde{W}}\left(D_{2}y\right)\bm{I}_{\left\{D_{3}y=\tilde{d}\right\}}(y),$
@@
[…]
@@
$=\frac{f_{Y\mid\tilde{V}=\tilde{d}}(y)\bm{I}_{\left\{y\;:\;y^{\top}D_{2}y-cy^{\top}D_{1}y\geq-\tilde{d}^{\top}\tilde{d}\right\}}(y)}{\Pr\left(\tilde{W}^{\top}\tilde{W}-c\tilde{Z}^{\top}\tilde{Z}\geq-\tilde{d}^{\top}\tilde{d}\right)}$
$=\frac{f_{\tilde{Z}}\left(D_{1}y\right)f_{\tilde{W}}\left(D_{2}y\right)\bm{I}_{\left\{D_{3}y=\tilde{d}\right\}}(y)\bm{I}_{\left\{y\;:\;y^{\top}D_{2}y-cy^{\top}D_{1}y\geq-\tilde{d}^{\top}\tilde{d}\right\}}(y)}{\Pr\left(\tilde{W}^{\top}\tilde{W}-c\tilde{Z}^{\top}\tilde{Z}\geq-\tilde{d}^{\top}\tilde{d}\right)}$,
@@
We now comment on Fig. B.1. When the overall null $H_{0}^{1:p}$ (2) holds and $H_{0}^{1:p}$ is rejected, the standard MLE has a much larger spread than the conditional MLE derived in this section, and indeed the mode of the former is not even at 0. This is because the overall null is rejected when the F-statistic is sufficiently large, which tends to occur when $\hat{\beta}_{1}$ and/or $\hat{\beta}_{2}$ is large. Thus, the standard MLE for $\beta_{1}$ will tend to be far from zero conditional on the rejection of the overall null. The selective estimate, on the other hand, adjusts for the selection event, and therefore has its mode at the true value 0 and is more tightly concentrated around it.
Figure B.1: We generate datasets $(X,Y)$ with $n=10$, $p=2$, and $\sigma^{2}=2$ under the overall null $H_{0}^{1:p}$ until we have $1\text{\times}{10}^{4}$ that satisfy (6). For each dataset, we compute the standard, unconditional MLE of $\beta$, $\hat{\beta}_{\text{MLE}}=(X^{\top}X)^{-1}X^{\top}Y$. For the same datasets, we compute the conditional MLE of $\beta_{1}$ that arises from numerically maximizing the conditional likelihood in (B.9). Each dataset corresponds to a point on the above scatterplots, with the standard point estimate on the horizontal axis and the conditional point estimate on the vertical axis. The point estimates in the left plot are computed using the true value of $\sigma^{2}$, and the point estimates in the right plot use the variance estimate $\tilde{\sigma}^{2}$} in (18).